OpenStudy (anonymous):

can you help me with this one ∫0Π/2cos(x)exdx i know i integrate twice so u= e^x dv= cos(x)dx du= e^x v= sin(x) u= e^x dv= sin(x) dx du= e^x v = -cos(x) ∫0Π/2cos(x)exdx=[(ex)(sin(x))−(ex)(−cos(x))]

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi/2} \cos(x) e^x dx\]

OpenStudy (anonymous):

this is how the problem is written

OpenStudy (mathmate):

Have you done integration by parts?

OpenStudy (anonymous):

I go the u = e^x dv= cos(x)dx du= e^x dx v= sin(x) then u =e^x du=e^x dv=sin(x) v= -cos(x)

OpenStudy (anonymous):

so... \[\int\limits_{0}^{\Pi/2} \cos(x) e^x dx= [e^x \sin(x)-e^x-(\cos(x))]\]

OpenStudy (anonymous):

of integral 0 to \[\Pi/2\] + integral 0 to \[\Pi/2\] e^x(-cos(x)) dx

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Let I=integral, and when you integrate by parts the first time, you get \(I=cos(x)e^x+\int sin(x)e^xdx\) when you integrate the second time, \(I=cos(x)e^x+sin(x)e^x-\int cos(x)e^xdx\) or \(I=cos(x)e^x+sin(x)e^x-I\) \( 2I=cos(x)e^x+sin(x)e^x\) hence \(\huge I=\frac{e^x(cos(x)+sin(x))}{2}\) I will leave it to you to evaluate the definite integral.

OpenStudy (anonymous):

is it 2.418040 ?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

I get a smaller number. Sure the limits are 0 to \(\pi/2\)? Want to check you work? Don't forget that e^0=1.